OpenStudy (anonymous):

Solve by substitution method Solution Set is? x+y=-9 y=-4x

OpenStudy (anonymous):

since you are given y=-4x, you can substitute( -4x) for y in the other eq and then you will have an eq in one variable, x, which you can then solve for

OpenStudy (anonymous):

yes i got that

OpenStudy (anonymous):

x+(-4x)=-9 x-4x=-9 -3x=-9 x=3

OpenStudy (anonymous):

so what did you get for x?

OpenStudy (anonymous):

annas has it.

OpenStudy (anonymous):

its a real simple question

OpenStudy (anonymous):

(-4x)+y=-9 x=3 my comp is slow :(

OpenStudy (anonymous):

but it says type in a orderd pair which is where i get lost

OpenStudy (anonymous):

(x,y) replace

OpenStudy (anonymous):

(3,-12)

OpenStudy (anonymous):

so you just multiply

OpenStudy (anonymous):

once you get the x value, plug it into either eq to get the y value. You then have the x and y value for the ordered pair

OpenStudy (anonymous):

so you would put like -4(3)=-12

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I put 3 and -12 in and it said wrong hmmm

OpenStudy (anonymous):

it worked i forgot the swiggless thingys